# VITA-Audio: Fast Interleaved Cross-Modal Token Generation for Efficient Large Speech-Language Model

### v1

@@ Abstract
With the growing requirement for natural human-computer interaction, speech-based systems receive increasing attention as speech is one of the most common forms of daily communication. However, the existing speech models still experience high latency when generating the first audio token during streaming, which poses a significant bottleneck for deployment. To address this issue, we propose VITA-Audio, an end-to-end large speech model with fast audio-text token generation. Specifically, we introduce a lightweight Multiple Cross-modal Token Prediction (MCTP) module that efficiently generates multiple audio tokens within a single model forward pass, which not only accelerates the inference but also significantly reduces the latency for generating the first audio in streaming scenarios. In addition, a four-stage progressive training strategy is explored to achieve model acceleration with minimal loss of speech quality. To our knowledge, VITA-Audio is the first multi-modal large language model capable of generating audio output during the first forward pass, enabling real-time conversational capabilities with minimal latency. VITA-Audio is fully reproducible and is trained on open-source data only. Experimental results demonstrate that our model achieves an inference speedup of 35×3$\sim$ 5$\times$3 ∼ 5 × at the 7B parameter scale, but also significantly outperforms open-source models of similar model size on multiple benchmarks for automatic speech recognition (ASR), text-to-speech (TTS), and spoken question answering (SQA) tasks.
@@ 1 Introduction
Speech interaction is becoming the primary modality for human-computer interaction due to its naturalness and its ability to convey richer informational dimensions through paralinguistic features such as prosody and intonation. Real-time speech systems have thus become a crucial research focus for enabling natural dialogue.
Traditional speech systems predominantly adopt a modular design that decomposes real-time speech processing into three discrete components: automatic speech recognition (ASR), large language models (LLMs), and text-to-speech (TTS) MThreads-full-duplex ; AudioGPT ; InternLM-XComposer2.5-OmniLive . However, this cascaded approach suffers from cumulative latency, loss of paralinguistic information (e.g., emotional prosody, rhythm) during modality conversion, and error accumulation between modules, substantially lowering the practical utility of cascaded architectures in real-time interactive scenarios. To address the limitations of traditional methods, many recent studies have adopted an end-to-end approach to handle inputs and outputs of the model MiniOmni ; SLAM-Omni ; LLaMA-Omni . These methods directly input speech into LLMs through an audio encoder and then synthesize speech response using the discrete tokens GLM-4-Voice or hidden states Freeze-Omni output by LLMs.
@@
Figure 1: (a) The audio sequence generated by the speech model exhibits a strong correlation with the corresponding text tokens. (b) With irrelevant text tokens being masked out, the model is still able to generate the correct audio, and the pronunciation remains contextually appropriate. However, if all text tokens are masked, the model outputs random audio. This suggests that the hidden states from the LLM include sufficient contextual information for generating the corresponding audio tokens. Consequently, the mapping from text hidden states to audio tokens can be accomplished using relatively simple modules, without the need for the extensive semantic modeling typically required by LLMs.
In Fig. 1(b), we show a Chinese sentence with a homograph “{CJK}UTF8gbsn行” as an example. The pronunciation of this character can be /xing/ or /hang/ depending on its context. Our speech model correctly decides the character’s pronunciation to be the former, given the hidden states of historical inputs. We then modify the model inference process by masking out all text hidden states, except the one corresponding to the token “{CJK}UTF8gbsn一行”(/yixing/) before generating its corresponding audio tokens. This modification prevents subsequently generated audio tokens from directly attending to other text tokens, although they can still attend to previously generated audio tokens. We find that the subsequently generated audio tokens accurately produce the sounds as “yixing”, which remains contextually appropriate. The same observation also holds for other non-homograph tokens. We therefore argue that the hidden states from the LLM include sufficient contextual information for generating its corresponding audio tokens, and attending to additional texts is unnecessary. Finally, we experiment with masking out all text tokens. This time, the generated audio fails to align with its text and sounds like random non-speech babbles even though the model has access to the previously generated audio tokens.
@@
In this paper, we introduce VITA-Audio, a lightweight framework that uses separate efficient modules, named Multiple Cross-modal Token Prediction (MCTP), to efficiently generate audio responses from text embeddings and LLM hidden states. This approach enables obtaining both text tokens and an audio chunk in a single LLM forward pass, achieving zero delay in audio tokens. A comparison of the delay of the first audio token is presented in Table 1 , where we define “audio token delay” as the number of additional LLM forward steps required to generate the first audio token after the first LLM forward pass. We distinguish this delay from “audio generation delay” which is the number of additional LLM forward passes to generate a meaningful and consistent chunk of audio. This distinction exists because speech tokenizers can handle varied lengths of context window. For example, the CNN-based SNAC decoder SNAC requires at least $3\times 7$ tokens to generate $0.083$ seconds of audio that is consistent with neighboring audio chunks at its boundaries. VITA-Audio has both zero audio token delay and zero audio generation delay.
@@
  • VITA-Audio achieves remarkable end-to-end inference acceleration by generating ten audio tokens in a single forward pass, resulting in 35×3$\sim$ 5$\times$3 ∼ 5 × speedup when implemented on a 7B LLM while preserving the ability of high-quality speech synthesis.
@@
Model Audio Token Delay Leveraging Hidden States End-to-End
Freeze-Omni  Freeze-Omni Text Length
MinMo  MinMo 5
Mini-Omni  MiniOmni 7
Moshi  Moshi 1
GLM-4-Voice  GLM-4-Voice 13
LUCY  LUCY 7
VITA-Audio 0
@@ 2 Related Work
Large language models have recently transformed human-computer interaction through their powerful natural language processing abilities. Extending their power to other modalities has since become the next step to explore. As one of the most natural forms of communication, speech interaction systems have thus become a major research focus. Traditional systems MThreads-full-duplex ; AudioGPT ; InternLM-XComposer2.5-OmniLive typically employ a cascaded architecture, a combination of separate ASR, LLMs, and TTS modules. However, this design presents several significant challenges, including accumulation of latency, loss of paralinguistic information, and error propagation across modules. Recent studies VITA ; Step-Audio have attempted to integrate audio encoders directly into LLMs using trainable adapters, which significantly improves system integration. However, these systems still rely on an independent TTS module to generate audio.
To enrich the input information of the speech decoder module, one research direction is to integrate the hidden states of LLMs into the speech decoder module. Llama-Omni LLaMA-Omni utilizes a non-autoregressive transformer to predict discrete audio tokens from upsampled hidden states and is optimized using the Connectionist Temporal Classification loss. Freeze-omni Freeze-Omni maximizes the retention of language capabilities by freezing the language model and employs a combination of autoregressive and non-autoregressive speech decoders, predicting audio tokens from the language model’s hidden states and text embeddings. Minmo MinMo integrates a language model with CosyVoice2 CosyVoice2 , where the language model processes pure text, and the CosyVoice2 mixed speech-text sequences.
End-to-end systems further integrate TTS functionality into LLMs, allowing the LLM to generate not only text responses but also speech responses. Currently, end-to-end speech models are primarily divided into two categories: parallel audio-text modeling and interleaved audio-text modeling. In the parallel modeling paradigm, the model uses different heads to process hidden states, generating both text and multiple audio tokens SLAM-Omni ; Moshi . Since the input to the LLM is altered during autoregression, maintaining the original capabilities of the LLM presents significant challenges. To perform inference without large-scale pretraining, Mini-Omni MiniOmni and LUCY LUCY rely on batch parallel decoding to preserve the inference capability of the LLM.
In the interleaved modeling paradigm GLM-4-Voice ; Spirit-LM , text and audio tokens are combined into a sequence, with the model alternating between predicting text and audio tokens. Compared to parallel-paradigm models, interleave-paradigm models appear to better preserve language capabilities, as suggested by the performance comparison on spoken question-answering benchmarks GLM-4-Voice . We attribute this difference to the fact that parallel-paradigm models use an average of text and audio representations as input, which significantly diverges from the inputs used during pretraining. However, interleave-paradigm models face a latency issue due to their sequential prediction of audio tokens, especially when the audio token rate is high.
@@ 3 Method
Figure 2: Architecture overview. (a) VITA-Audio is an end-to-end large speech model equipped with 10 light-weight Multiple Cross-modal Token Prediction (MCTP) modules that enable speech generation with extremely low latency. As shown in Fig. 1, we observe that the hidden states of certain text tokens in the LLM backbone contain sufficient semantic information for generating the corresponding audio tokens, which means that it is unnecessary to attend to additional text tokens when generating audio. Thus, we propose to utilize a set of light-weight MCTP modules to model the mapping from LLM hidden states to the audio tokens. (b) The details of the MCTP modules. Our MCTP module has a light-weight architecture, which enables it to finish one forward pass within $0.0024$ seconds ( $11$% of the LLM backbone). The MCTP module is capable of generating 10 audio tokens from the LLM hidden states and the text embedding, and the generated audio tokens can be decoded by the audio decoder directly. The utilization of MCTP modules enables VITA-Audio to generate audio responses in one LLM forward pass, which achieves extremely fast generation speed.
@@ 3.1 Overview
Table 2: Summary of datasets used in VITA-Audio for different stages.
Figure 4: The four text-audio interleaved inference modes are illustrated as follows: 1) Turbo: As the fastest inference mode, it generates $1$ token by the main model and $10$ additional tokens via MCTP in each forward pass. To ensure that a valid audio chunk is decoded after the first forward pass, the first generated $11$ tokens are split into $1$ text token and $10$ audio tokens. Then, the Turbo mode iteratively generates $4$ text tokens and $10$ audio tokens in the following forward. 2) Boost: To enhance the quality of text tokens, Boost mode follows the text-audio cyclic pattern of Turbo mode, with the main model generating every text token and MCTP generating every audio token. 3) Balance: To keep a balanced text-audio ratio, i.e., $1:2$, the balance mode further changes the text-audio cyclic pattern of the Boost mode. Specifically, the balance mode sequentially generates $1$ text token from the main model, $4$ audio tokens ($2$ tag tokens mark the beginning and end of audios, and $2$ common tokens denote the audio content) from MCTP, $3$ text tokens from the main model, $8$ text tokens ($2$ tag tokens mark the beginning and end of audios, and $6$ common tokens denote the audio content) from MCTP, and then iteratively generates $4$ text tokens from the main model and $10$ audio tokens ($2$ tag tokens mark the beginning and end of audios, and $8$ common tokens denote the audio content) from MCTP. 4) Vanilla: As the slowest inference mode, Vanilla mode follows the text-audio cyclic pattern of Balance mode, with the main model generating every token.
As illustrated in Fig. 2, VITA-Audio consists of four major components: an audio encoder, an audio decoder, a large language model backbone, and a set of Cross-modal Token Prediction (MCTP) modules. Similarly to GLM-4-Voice GLM-4-Voice , we use CosyVoice CosyVoice as the audio encoder and decoder. The audio signal is first encoded into a sequence of discrete audio tokens through the audio encoder, which are subsequently fed into the LLM for processing. During each forward pass, the LLM alternately generates text and audio tokens. The hidden states from the final layer of LLM, along with the embedding of the predicted token, are provided as input to the MCTP modules. The historical input tokens, the tokens predicted by the LLM, and by the MCTP modules are concatenated to form the inputs to the next LLM forward pass. Finally, the audio tokens generated by both the LLM and the MCTP modules are aggregated and passed to the audio decoder to generate the final audio output.
@@ 3.3.1 Data Construction
VITA-Audio is trained exclusively on open-source datasets, integrating multi-domain and multi-language speech data resources. The training dataset encompasses a diverse range of sources. Detailed descriptions of the datasets and the data proportions used at each stage are provided in Table 2.
@@ 3.3.2 Training Pipline
For VITA-Audio to output a consistent sequence of audio tokens in a single forward pass, each MCTP module needs to model a different distribution. As a result, training all the MCTP modules simultaneously becomes a challenging task, especially when the number of modules is large, because their optimization objectives are not guaranteed to align. We propose a four-stage training strategy, as illustrated in Fig. 3, to progressively equip the MCTP modules with the ability to map text to its audio, thereby reducing the difficulty of their convergence.
Stage 4: Supervised Fine-tuning. After the previous three training stages, VITA-Audio has acquired the ability to efficiently and accurately map text to audio. To enable speech-to-speech dialogue capability, we then conduct supervised fine-tuning using speech QA datasets while maintaining a small amount of TTS, ASR, and text-only data to ensure training stability. To optimize training effectiveness, different learning rates are used for the MCTP module and the main LLM. For the speech-to-speech data, we employ an interleaved output format. This design enforces the model to initiate audio token generation during the first forward pass, enabling synchronized decoding of the audio tokens rather than waiting until all text tokens have been generated. This interleaved data format is exemplified as follows:
Speech QA Interleaved Data Format
    "messages": [
        "role": "user",
        "content": "<|begin_of_audio|> audio_sequence <|end_of_audio|>"
        "role": "assistant",
        "content": "text_sequence_1 <|begin_of_audio|>
        audio_sequence_1 <|end_of_audio|> text_sequence_2
        <|begin_of_audio|> audio_sequence_2 <|end_of_audio|>"
@@ 3.4 Inference
As shown in Fig. 4, we have designed four different inference paradigms to address various scenarios.
@@
For speech dialogue tasks, we introduce VITA-Audio-Speed and VITA-Audio-Balance. Their main difference lies in audio token generation: VITA-Audio-Speed generates eight directly decodable audio tokens in the first forward pass, whereas VITA-Audio-Balance adheres to a strict $1:2$ text-to-audio token ratio for enhanced speech quality. The latter requires two forward passes to generate enough audio tokens for decoding. To optimize model performance, distinct models were trained for each of the two modes.
@@ 4.1 Experiment Settings
We use the pretrained Qwen2.5-7B Qwen2.5 as the pre-trained text LLM. The initial version of VITA-Audio utilizes the speech tokenizer and speech decoder in GLM-4-Voice GLM-4-Voice , which effectively captures semantic information at an ultra-low bitrate. In the second version, i.e., VITA-Audio-Plus further replaces the GLM-4-Voice tokenizer with SenseVoiceSmall FunAudioLLM and an MLP-based adapter. The detail comparison between VITA-Audio and VITA-Audio-Plus is listed in Table 3.
Throughout the training process, the global batch size is set to $256$. In Stage 1, the learning rate for the LLM is $6e^{-5}$. In Stages 2 and 3, the learning rate for the MCTP is $1e{-3}$. In Stage 4, the learning rate for the LLM is $5e{-5}$, while the learning rate for the MCTP is $4e{-5}$. The model is trained for $1$ epoch. We use the AdamW AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. We evaluate VITA-Audio ’s performance on ASR, TTS, and SQA tasks, respectively.
Table 3: Comparison of model structures between VITA-Audio and VITA-Audio-Plus.
[…]
@@
Model #Params Llama Question TriviaQA Web Question Mean S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S Proprietary Models MinMo  MinMo 7B 78.9 64.1 48.3 37.5 55.0 39.9 60.7 47.2 Open-source Models Moshi  Moshi 7B 62.3 21.0 22.8 7.3 26.6 9.2 37.2 12.5 GLM-4-Voice  GLM-4-Voice 9B 64.7 50.7 39.1 26.5 55.0 39.9 45.3 31.0 LUCY (S2)  LUCY 7B 59.6 51.0 23.2 18.2 26.6 18.2 36.5 29.1 VITA-Audio-Boost 7B 68.7 60.3 30.5 29.3 32.9 30.4 44.0 40.0 VITA-Audio-Vanilla 7B 71.3 66.3 31.9 30.1 33.5 31.4 45.6 42.6 VITA-Audio-Plus-Vanilla 7B 75.6 68.0 45.9 42.7 45.0 41.7 55.5 50.8
[…]
@@
Model Seed-TTS LibriTTS test-zh test-en test-hard test-clean CER (%) $\downarrow$ WER (%) $\downarrow$ WER (%) $\downarrow$ WER (%) $\downarrow$ Seed-TTS  Seed-TTS 1.12 2.25 7.59 CosyVoice  CosyVoice 3.63 4.29 11.75 2.89 CosyVoice2  CosyVoice2 1.45 2.57 6.83 2.47 VITA-1.5 (S3)  VITA-1.5 8.44 2.63 GLM-4-Voice  GLM-4-Voice 2.91 2.10 5.64 VITA-Audio-Turbo (S1) 1.18 1.92 10.58 1.96 VITA-Audio-Turbo (S2) 0.96 1.92 9.72 1.98 VITA-Audio-Turbo (S3) 1.05 1.77 9.86 1.99 VITA-Audio-Turbo 1.07 2.26 10.08 2.08 VITA-Audio-Plus-Vanilla 1.13 1.85 10.21 1.89
[…]
Datasets Model WER (%) $\downarrow$ LibriSpeech Librispeech test-clean | test-other Qwen2-Audio-base  Qwen2-Audio 1.74 | 4.04 Baichuan-Audio-base  Baichuan-Audio 3.02 | 6.04 Freeze-Omni  Freeze-Omni 3.82 | 9.79 VITA-1.5 (S3)  VITA-1.5 8.14 | 18.41 LUCY (S1)  LUCY 3.36 | 8.05 Step-Audio-chat  Step-Audio 3.19 | 10.67 Qwen2.5-Omni  Qwen2.5-Omni 2.37 | 4.21 VITA-Audio-Turbo 6.29 | 12.86 VITA-Audio-Vanilla 2.98 | 8.07 VITA-Audio-Plus-Vanilla (S1) 1.91 | 4.29 VITA-Audio-Plus-Vanilla 2.00 | 4.60 Fleurs Fleurs zh | en Qwen2-Audio-base  Qwen2-Audio 3.63 | 5.20 Baichuan-Audio-base  Baichuan-Audio 4.15 | 8.07 Step-Audio-chat  Step-Audio 4.26 | 8.56 Qwen2.5-Omni  Qwen2.5-Omni 2.92 | 4.17 VITA-Audio-Plus-Vanilla 3.69 | 4.54
[…]
@@
Datasets Model WER (%) $\downarrow$ AISHELL-1 AISHELL_1 Qwen2-Audio-base  Qwen2-Audio 1.52 Baichuan-Audio-base  Baichuan-Audio 1.93 Freeze-Omni  Freeze-Omni 2.48 LUCY (S1)  LUCY 2.40 Step-Audio-chat  Step-Audio 2.14 Qwen2.5-Omni  Qwen2.5-Omni 1.13 VITA-Audio-Turbo 7.70 VITA-Audio-Vanilla 4.46 VITA-Audio-Plus-Vanilla (S1) 1.51 VITA-Audio-Plus-Vanilla 1.94 AISHELL-2 ios AISHELL_2 Qwen2-Audio-base  Qwen2-Audio 3.08 Baichuan-Audio-base  Baichuan-Audio 3.87 Step-Audio-chat  Step-Audio 3.89 Qwen2.5-Omni  Qwen2.5-Omni 2.56 VITA-Audio-Plus-Vanilla 3.29 WenetSpeech WenetSpeech test-meeting | test-net Qwen2-Audio-base  Qwen2-Audio 8.40 | 7.64 Baichuan-Audio-base  Baichuan-Audio 13.28 | 10.13 Freeze-Omni  Freeze-Omni 11.80 | 13.46 VITA-1.5 (S3)  VITA-1.5 8.14 | 18.41 LUCY (S1)  LUCY 10.42 | 10.42 Step-Audio-chat  Step-Audio 10.83 | 9.47 Qwen2.5-Omni  Qwen2.5-Omni 7.71 | 6.04 VITA-Audio-Turbo 18.66 | 23.97 VITA-Audio-Vanilla 13.45 | 17.34 VITA-Audio-Plus-Vanilla (S1) 6.68 | 6.59 VITA-Audio-Plus-Vanilla 7.12 | 6.90

### v2

@@ Abstract
With the growing requirement for natural human-computer interaction, speech-based systems receive increasing attention as speech is one of the most common forms of daily communication. However, the existing speech models still experience high latency when generating the first audio token during streaming, which poses a significant bottleneck for deployment. To address this issue, we propose VITA-Audio, an end-to-end large speech model with fast audio-text token generation. Specifically, we introduce a lightweight Multiple Cross-modal Token Prediction (MCTP) module that efficiently generates multiple audio tokens within a single model forward pass, which not only accelerates the inference but also significantly reduces the latency for generating the first audio in streaming scenarios. In addition, a four-stage progressive training strategy is explored to achieve model acceleration with minimal loss of speech quality. To our knowledge, VITA-Audio is the first multi-modal large language model capable of generating audio output during the first forward pass, enabling real-time conversational capabilities with minimal latency. VITA-Audio is fully reproducible and is trained on open-source data only. Experimental results demonstrate that our model achieves an inference speedup of $3\sim 5\times$ at 7B parameter scale, but also significantly outperforms open-source models of similar model size on multiple benchmarks for automatic speech recognition (ASR), text-to-speech (TTS), and spoken question answering (SQA).
@@ 1 Introduction
Real-time speech systems have become a crucial research focus for enabling natural dialogue. Traditional speech systems predominantly adopt a modular design that decomposes real-time speech processing into three discrete components: automatic speech recognition (ASR), large language models (LLMs), and text-to-speech (TTS) MThreads-full-duplex ; AudioGPT ; InternLM-XComposer2.5-OmniLive . However, this cascaded approach suffers from cumulative latency, loss of paralinguistic information (e.g., emotional prosody, rhythm) during modality conversion, and error accumulation between modules, substantially lowering the practical utility of cascaded architectures in real-time interactive scenarios. To address the limitations of traditional methods, many recent studies have adopted an end-to-end approach to handle inputs and outputs of the model MiniOmni ; SLAM-Omni ; LLaMA-Omni . These methods directly input speech into LLMs through an audio encoder and then synthesize speech response with discrete tokens GLM-4-Voice or LLM hidden states Freeze-Omni .
@@
Figure 1: (a) The audio sequence generated by the speech model exhibits a strong correlation with the corresponding text tokens. (b) With irrelevant text tokens being masked out, the model is still able to generate the correct audio, and the pronunciation remains contextually appropriate. However, if all text tokens are masked, the model outputs random audio. This suggests that the hidden states from the LLM include sufficient contextual information for generating the corresponding audio tokens. Consequently, the mapping from text hidden states to audio tokens is accomplished using relatively simple modules, without the need for the extensive semantic modeling typically required by LLMs.
In Fig. 1(b), we show a Chinese sentence with a homograph “行” as an example. The pronunciation of this character can be /xing/ or /hang/ depending on its context. In Chinese, ’yihang’ emphasizes spatial arrangement (e.g. a row/line of egrets), while ’yixing’ focuses on the concept of a group or unit (e.g. a group/party of people). Our speech model correctly decides the character’s pronunciation to be the former, given the hidden states of historical inputs. We then modify the model inference process by masking out all text hidden states, except the one corresponding to the token “一行”(/yixing/) before generating its corresponding audio tokens. This modification prevents subsequently generated audio tokens from directly attending to other text tokens, although they can still attend to previously generated audio tokens. We find that the subsequently generated audio tokens accurately produce the sounds as “yixing”, which remains contextually appropriate. The same observation also holds for other non-homograph tokens. We therefore argue that the hidden states from the LLM include sufficient contextual information for generating its corresponding audio tokens, and attending to additional texts is unnecessary. Finally, we experiment with masking out all text tokens. This time, the generated audio fails to align with its text and sounds like random non-speech babbles even though the model has access to the previously generated audio tokens.
@@
In this paper, we introduce VITA-Audio, a lightweight framework that uses separate efficient modules, named Multiple Cross-modal Token Prediction (MCTP), to efficiently generate audio responses from text embeddings and LLM hidden states. This approach enables obtaining both text tokens and an audio chunk in a single LLM forward pass, achieving zero delay in audio tokens. A comparison of the delay of the first audio token is presented in Table 1 , where we define “audio token delay” as the number of additional LLM forward steps required to generate the first audio token after the first LLM forward pass. We distinguish this delay from “audio generation delay” which is the number of additional LLM forward passes to generate a meaningful and consistent chunk of audio.
VITA-Audio has both zero audio token delay and zero audio generation delay.
@@
  • VITA-Audio achieves remarkable end-to-end inference acceleration by generating ten audio tokens in a single forward pass, resulting in $3\sim 5\times$ speedup when implemented on a 7B LLM while preserving the ability of high-quality speech synthesis.
@@
Model Audio Token Delay Leveraging Hidden States End-to-End Freeze-Omni  Freeze-Omni Text Length MinMo  MinMo 5 Mini-Omni  MiniOmni 7 Moshi  Moshi 1 GLM-4-Voice  GLM-4-Voice 13 LUCY  LUCY 7 VITA-Audio 0
@@ 2 Related Work
Large language models (LLMs) have revolutionized human-computer interaction with advanced natural language processing. Extending these capabilities to speech—a natural communication modality—has become a key research focus. Traditional speech interaction systems MThreads-full-duplex ; AudioGPT ; InternLM-XComposer2.5-OmniLive typically adopt a cascaded architecture, combining separate ASR, LLM, and TTS modules. However, this approach suffers from increased latency, loss of paralinguistic cues, and error propagation.
Recent work VITA ; Step-Audio has improved integration by connecting audio encoders to LLMs via trainable adapters, but still relies on independent TTS modules. To address this, some methods incorporate LLM hidden states into audio decoders. For example, Llama-Omni LLaMA-Omni uses a non-autoregressive transformer to predict audio tokens from upsampled LLM states, while Freeze-omni Freeze-Omni freezes the LLM and combines autoregressive and non-autoregressive decoders. Minmo MinMo integrates a language model with CosyVoice2 CosyVoice2 for mixed speech-text processing.
End-to-end models further unify TTS within LLMs, enabling direct text and speech generation. These models follow either parallel or interleaved audio-text modeling. In the parallel modeling paradigm, the model uses different heads to process hidden states, generating both text and multiple audio tokens SLAM-Omni ; Moshi . Since the input to the LLM is altered during autoregression, maintaining the original capabilities of the LLM presents significant challenges. To perform inference without large-scale pretraining, Mini-Omni MiniOmni and LUCY LUCY rely on batch parallel decoding to preserve the inference capability of the LLM.
Compared to parallel-paradigm models, interleave-paradigm models appear to better preserve language capabilities, as suggested by the performance comparison on spoken question-answering benchmarks GLM-4-Voice . We attribute this difference to the fact that parallel-paradigm models use an average of text and audio representations as input, which significantly diverges from the inputs used during pretraining. However, interleave-paradigm models face a latency issue due to their sequential prediction of audio tokens, especially when the audio token rate is high.
@@ 3 Method
Figure 2: Architecture overview. (a) VITA-Audio is an end-to-end large speech model equipped with 10 light-weight Multiple Cross-modal Token Prediction (MCTP) modules that enable speech generation with extremely low latency. As shown in Fig. 1, we observe that the hidden states of certain text tokens in the LLM backbone contain sufficient semantic information for generating the corresponding audio tokens, which means that it is unnecessary to attend to additional text tokens when generating audio. Thus, we propose to utilize a set of light-weight MCTP modules to model the mapping from LLM hidden states to the audio tokens. (b) The details of the MCTP modules. Our MCTP module has a light-weight architecture, which enables it to finish one forward pass within $0.0024$ seconds ( 1111% of the LLM backbone). The MCTP module is capable of generating 10 audio tokens from the LLM hidden states and the text embedding, and the generated audio tokens can be decoded by the audio decoder directly. The utilization of MCTP modules enables VITA-Audio to generate audio responses in one LLM forward pass, which achieves extremely fast generation speed.
@@ 3.1 Overview
As illustrated in Fig. 2, VITA-Audio consists of four major components: an audio encoder, an audio decoder, a large language model backbone, and a set of Cross-modal Token Prediction (MCTP) modules. The audio signal is first processed by the audio encoder, whose output is then fed into the LLM for further processing. During each forward pass, the LLM alternately generates text and audio tokens. The hidden states from the final layer of LLM, along with the embedding of the predicted token, are provided as input to the MCTP modules. The historical input tokens, the tokens predicted by the LLM, and by the MCTP modules are concatenated to form the inputs to the next LLM forward pass. Finally, the audio tokens generated by both the LLM and the MCTP modules are aggregated and passed to the audio decoder to generate the final audio output.
@@ 3.3.1 Data Construction
VITA-Audio is trained exclusively on open-source datasets, integrating multi-domain and multi-language speech data resources. The training dataset encompasses a diverse range of sources. Detailed descriptions of the datasets used at each stage are provided in Table E1 in the Appendix.
@@ 3.3.2 Training Pipline
For VITA-Audio to output a consistent sequence of audio tokens in a single forward pass, each MCTP module must model a distinct distribution. As a result, training all the MCTP modules simultaneously becomes a challenging task, especially when the number of modules is large, due to potentially misaligned optimization objectives. We propose a four-stage training strategy, as shown in Fig. 3, to progressively equip the MCTP modules with the ability to map text to its audio, thereby reducing the difficulty of their convergence. Further training details are provided in Sec. B of the Appendix.
Figure 4: The four text-audio interleaved inference modes are illustrated as follows: 1) Turbo: As the fastest inference mode, it generates 11 token by the main model and $10$ additional tokens via MCTP in each forward pass. To ensure that a valid audio chunk is decoded after the first forward pass, the first generated $11$ tokens are split into $1$ text token and $10$ audio tokens. Then, the Turbo mode iteratively generates $4$ text tokens and $10$ audio tokens in the following forward. 2) Boost: To enhance the quality of text tokens, Boost mode follows the text-audio cyclic pattern of Turbo mode, with the main model generating every text token and MCTP generating every audio token. 3) Balance: To keep a balanced text-audio ratio, i.e., $1:2$, the balance mode further changes the text-audio cyclic pattern of the Boost mode. Specifically, the balance mode sequentially generates $1$ text token from the main model, 44 audio tokens (22 tag tokens mark the beginning and end of audios, and 22 common tokens denote the audio content) from MCTP, 33 text tokens from the main model, 88 text tokens (22 tag tokens mark the beginning and end of audios, and 66 common tokens denote the audio content) from MCTP, and then iteratively generates $4$ text tokens from the main model and $10$ audio tokens (22 tag tokens mark the beginning and end of audios, and 88 common tokens denote the audio content) from MCTP. 4) Vanilla: As the slowest inference mode, Vanilla mode follows the text-audio cyclic pattern of Balance mode, with the main model generating every token.
@@ 3.4 Inference
In order to address diverse scenarios, four distinct inference paradigms have been designed as shown in Fig. 4.
@@
For speech dialogue tasks, we introduce VITA-Audio-Boost and VITA-Audio-Balance. Their main difference lies in audio token generation: VITA-Audio-Boost generates eight directly decodable audio tokens in the first forward pass, whereas VITA-Audio-Balance adheres to a strict $1:2$ text-to-audio token ratio for enhanced speech quality. The latter requires two forward passes to generate enough audio tokens for decoding. To optimize model performance, distinct models were trained for each of the two modes.
@@ 4.1 Experiment Settings
We use the Qwen2.5-7B-Instruct Qwen2.5 as the pre-trained text LLM. The initial version of VITA-Audio utilizes the speech tokenizer and speech decoder in GLM-4-Voice GLM-4-Voice , which effectively captures semantic information at an ultra-low bitrate. In the second version, i.e., VITA-Audio-Plus further replaces the GLM-4-Voice tokenizer with SenseVoiceSmall FunAudioLLM and an MLP-based adapter. The detail comparison between VITA-Audio and VITA-Audio-Plus is listed in Table E2.
Table 2: Results on Spoken Question Answering (SQA) benchmarks. “Sxx” denotes the $x$-th training stage of speech models.
@@
Model #Params Llama Question TriviaQA Web Question Mean S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S Proprietary Models MinMo  MinMo 7B 78.9 64.1 48.3 37.5 55.0 39.9 60.7 47.2 Open-source Models Moshi  Moshi 7B 62.3 21.0 22.8 7.3 26.6 9.2 37.2 12.5 GLM-4-Voice  GLM-4-Voice 9B 64.7 50.7 39.1 26.5 55.0 39.9 45.3 31.0 LUCY (S2)  LUCY 7B 59.6 51.0 23.2 18.2 26.6 18.2 36.5 29.1 VITA-Audio-Boost 7B 68.7 60.3 30.5 29.3 32.9 30.4 44.0 40.0 VITA-Audio-Vanilla 7B 71.3 66.3 31.9 30.1 33.5 31.4 45.6 42.6 VITA-Audio-Plus-Boost 7B 76.3 64.6 43.6 39.5 44.2 40.0 54.7 48.0 VITA-Audio-Plus-Vanilla 7B 75.6 68.0 45.9 42.7 45.0 41.7 55.5 50.8
Table 3: Results on Text to Speech (TTS) Benchmarks. “Sxx” denotes the $x$-th training stage.
@@
Model Seed-TTS LibriTTS test-zh test-en test-hard test-clean CER (%) $\downarrow$ WER (%) $\downarrow$ WER (%) $\downarrow$ WER (%) $\downarrow$ Seed-TTS  Seed-TTS 1.12 2.25 7.59 CosyVoice  CosyVoice 3.63 4.29 11.75 2.89 CosyVoice2  CosyVoice2 1.45 2.57 6.83 2.47 VITA-1.5 (S3)  VITA-1.5 8.44 2.63 GLM-4-Voice  GLM-4-Voice 2.91 2.10 5.64 VITA-Audio-Turbo (S1) 1.18 1.92 10.58 1.96 VITA-Audio-Turbo (S2) 0.96 1.92 9.72 1.98 VITA-Audio-Turbo (S3) 1.05 1.77 9.86 1.99 VITA-Audio-Turbo (S4) 1.07 2.26 10.08 2.08 VITA-Audio-Plus-Boost 1.32 2.21 12.05 2.21 VITA-Audio-Plus-Vanilla 1.13 1.85 10.21 1.89
Table 4: Results on Automatic Speech Recognition (ASR) Benchmarks. “Sxx” denotes the $x$-th training stage. Compared to other methods, VITA-Audio is trained with open-source data only.
@@
Model WenetSpeech AIShell LibriSpeech test_meeting $\downarrow$ test_net $\downarrow$ test $\downarrow$ test-clean $\downarrow$ test-other $\downarrow$ Qwen2-Audio-base  Qwen2-Audio 8.40 7.64 1.52 1.74 4.04 Baichuan-Audio-base  Baichuan-Audio 13.28 10.13 1.93 3.02 6.04 VITA-1.5 (S3)  VITA 10.0 8.4 2.2 3.4 7.5 Freeze-Omni  Freeze-Omni 13.46 11.8 2.48 3.82 9.79 LUCY (S1)  LUCY 10.42 8.78 2.40 3.36 8.05 Step-Audio-chat  Step-Audio 10.83 9.47 2.14 3.19 10.67 Qwen2.5-Omni  Qwen2.5-Omni 7.71 6.04 1.13 2.37 4.21 VITA-Audio-Vanilla 17.34 13.45 4.46 2.98 8.07 VITA-Audio-Plus-Boost 9.38 8.97 4.72 3.13 7.07 VITA-Audio-Plus-Vanilla(S1) 6.68 6.59 1.51 1.91 4.29 VITA-Audio-Plus-Vanilla 7.12 6.90 1.94 2.00 4.60